OpenStudy (anonymous):

[4_02]Identify the correct x intercept and y intercept of the equation 2x - 3y = -18.

OpenStudy (anonymous):

x intercept = (9, 0); y intercept = (0, -6) x intercept = (6, 0); y intercept = (0, -9) x intercept = (-9, 0); y intercept = (0, 6) x intercept = (-6, 0); y intercept = (0, 9)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yeh i know i have done a mistake

OpenStudy (anonymous):

y intercept is 6

OpenStudy (anonymous):

change in the terms of y ie y=mx+c y=(2x+18)/3 hence y intercept is 6 now find for x

OpenStudy (anonymous):

x intercept is -9